OpenStudy (anonymous):

Write the given function in vertex form by completing the square: y=x^2+9x+2

myininaya (myininaya):

Ok! So \[x^2+9x+k\] How do you find k so that this is a complete square?

myininaya (myininaya):

\[x^2+9x+(\frac{9}{2})^2=(x+\frac{9}{2})^2\] This will always work if the coefficient of x^2 is 1 Just divide the number in front of x by 2 then square like I did above

myininaya (myininaya):

Now back to finding the vertex form!

myininaya (myininaya):

\[y=x^2+9x+(\frac{9}{2})^2-(\frac{9}{2})^2+2\] Whatever I add in I must also take out

myininaya (myininaya):

The form we are aiming for by the way is: \[y=a(x-h)^2+k\]

myininaya (myininaya):

So You should be able to write that one part (x^2+9x+(9/2))^2) as a complete square which I at the beginning of this thread

myininaya (myininaya):

And then all you have to do after simplify if you can such as the -(9/2)^2+2 part

OpenStudy (anonymous):

thankyou!